# Federated Semi-Supervised and Semi-Asynchronous Learning for Anomaly Detection in IoT Networks

### v2

@@ Abstract
Index Terms:
Anomaly Detection, Federated Learning, Internet of Things, Semi-Supervised Learning, Semi-Asynchronous Mechanism.
publicationid: pubid: $0$–$00$ 2021 IEEE
@@ I Introduction
Over the past few decades, with the development of sensors and wireless communication technologies, Internet of Things (IoT) networks have become a popular architecture to support many modern applications or services, such as smart homes, smart healthcare, smart cities, smart grid, and smart transportation, among many others [1]. However, simultaneously with the massive deployment of IoT networks, the attack surface where IoT devices may be hacked and exploited has grown in recent years, such as DDoS attack, Side-channel attack or Bot attack [2], which seriously hinders the development of IoT networks [3]. For this reason, effective countermeasures need to be developed to detect infected IoT devices in networks, in order to strengthen and ensure the security of IoT networks.
In recent years, artificial intelligence (AI) technologies, such as machine learning (ML) and deep learning (DL), have been widely used for attack detection in IoT networks due to their superior performance [4]. Depending on the detection location, traditional AI-based schemes can be categorized into centralized and distributed detection [5]. In centralized detection, IoT devices send their local data to a central server for detection. This inevitably introduces expensive communication costs, serious privacy leakage, and poor scalability, although it can achieve high detection accuracy. Distributed detection, known also as on-device detection, is performed locally on devices based on their private data. Nonetheless, it is limited by the lack of interaction with and profit from other devices, which leads to the problem of isolated data islands, resulting in low detection accuracy [6].
To address the aforementioned limitations, Federated Learning (FL) [7] comes into being, in which multiple devices (i.e., clients) cooperatively train a global model under the coordination of a central entity (i.e., the server) without the leakage of local data. At each round of FL, clients perform the local training and upload the model parameters instead of raw data, and then the server is responsible for aggregating and distributing the parameters to share the knowledge in the data of various clients, thus achieving a trade-off between privacy, accuracy, communication cost and latency. The FL procedure continues until the model convergence or after a specified number of rounds. Many FL-based approaches for anomaly detection in IoT networks have been proposed, and Table I summarizes some classic and state-of-the-art works [8, 9, 10, 11, 12]. Based on the comprehensive analysis, we find that the existing researches are still subject to the following challenges:
@@
Nguyen et al. [8] Full No Benign Synchronous None
Rey et al. [9] Full Yes Yes/Benign Synchronous None
Campos et al. [10] Full Yes Yes Synchronous None
Liu et al. [11] Full No Yes Synchronous Gradient Compression
Tian et al. [12] One Yes Benign Asynchronous Gradient Compensation
@@
  • Lack of Labeled Data: A common limitation of existing works is that they only consider the supervised learning scenarios, where the local data is fully annotated with ground truths. It is not realistic because users and end devices do not have the ability and corresponding expertise to label the data [13]. Furthermore, some works [8, 12] argue their schemes are unsupervised, but in fact, they are based on the assumption that all data in the model training phase is benign, which is essentially supervised. At the same time, this assumption is unreasonable because we can not guarantee it in real networks, and once the devices are compromised during the training phase, the malicious samples will seriously degrade the performance [14].
  • Device Heterogeneity: Although synchronous FL [8, 15, 16] is a natural choice, due to the vast differences in computing capacities, communication resources and data volumes, the required time for different devices to perform local updates and receive/upload models varies greatly [17]. Therefore, due to the device heterogeneity, the server has to wait for the slowest client in synchronous FL, which leads to unnecessary waiting time and low round efficiency. Although asynchronous FL can deal with the above problems, it requires frequent model updates, resulting in an expensive consumption of communication and computing resources. The server receives numerous local updates sent by clients, which can overwhelm the server and provide little benefit to the model convergence [18].
  • Non-IID Data: In IoT networks, some devices may function normally, while others may suffer from multiple attacks. As a result, the local data of devices can not be regarded as a uniform distribution of the overall sample. In other words, the data of different clients is often not independent-and-identically-distributed (Non-IID) [19]. It has been proved that in the presence of non-IID data, the performance of FL degrades significantly in terms of the convergence rate and model accuracy [20].
  • Significant Communication Cost: Although FL does not need to upload local raw data to the server, frequent updates of model parameters (up to tens or even hundreds of rounds) are still a huge communication overhead for resource-constrained IoT networks. Meanwhile, the server may become a communication performance bottleneck since it needs to frequently distribute/receive the models to/from clients [21].
@@
  • We implement extensive experiments on the CIC-IDS 2017 [22] dataset. The experimental results show that the performance of FedS3A is far superior to the state-of-the-art detection methods [8, 9, 10, 11, 12] based on federated learning [7, 23] in terms of the model performance, communication cost and round efficiency. It can achieve greater than 98% accuracy even when the data is non-IID and reduce higher than 50% communication cost. Furthermore, we conduct ablation experiments on FedS3A and evaluate the impact of different hyperparameters and functions to demonstrate its efficiency and effectiveness.
@@ II-A Federated Learning for Anomaly Detection in IoT Networks
Federated Learning has been widely used for anomaly detection in IoT networks due to cooperative learning and privacy preservation. DIoT [8] is the first attempt to apply the federated learning approach for anomaly detection in IoT networks. It assumes that IoT devices are not compromised during the model training phase, leaving enough time to utilize the Gate Recurrent Unit (GRU) model to learn benign behavior patterns. Abnormalities are detected by deviations from benign behaviors. Liu et al. [11] use Convolutional Neural Networks (CNN) to extract fine-grained features, and then the trained Long Short-Term Memory (LSTM) model is utilized for anomaly detection. In addition, in order to improve communication efficiency, a top-k selection-based gradient compression scheme is proposed to reduce the communication cost.
Rey et al. [9] studies the performance of federated learning in two scenarios based on the N-BaIoT [24] dataset. One is that the client-side data is all labeled, and MultiLayer Perceptron (MLP) is used to perform supervised learning for classification, while the other is that clients only have benign data and utilize AutoEncoder (AE) for anomaly detection training. Based on the CIC-ToN-IoT [25] dataset, Campos et al. [10] define three scenarios, namely basic, balanced, and mixed scenarios, according to the distribution of the client-side dataset. Meanwhile, a novel aggregation function called Fed+ is proposed, which is more suitable for non-IID and skewed training data, and Multiple Linear Regression (MLR) is used for model training. Tian et al. [12] focus on the gradient divergence and delay problem caused by the IoT device heterogeneity and non-IID data in federated learning. A method based on Taylor expansion is proposed to compensate for the inconsistency.
@@ II-B Federated Semi-Supervised Learning
Semi-supervised learning [26] has attracted increasing attention in the past few years. It aims to learn a model under massive unlabeled data with a little supervised information. Therefore, in the context of federated learning, a novel and promising paradigm called federated semi-supervised learning (FSSL) is proposed and quickly becomes a research hotspot [27]. There are two essential scenarios according to the location of the labeled data. The first is the conventional scenario where clients have both labeled and unlabeled data (labels-at-client), and the second is the disjoint scenario where the labeled data is only available at the server (labels-at-server) [13].
Zhu et al. [28] combine the CNN-GRU model with the pseudo-labeling technique for FSSL to identify the travel model. Meanwhile, a group-based aggregation function and a flip-based data augmentation (DA) scheme are proposed to improve the performance and convergence speed of the trained model. Nandury et al. [29] analyze the influence of the update diversity of model parameters on model convergence, and propose diversity scaling to compensate for the effect of inter-client update dissimilarity. Zhang et al. [30] propose a metric to measure the non-IID distribution of inter-client data. Pseudo-labels generated by weak DA samples are used to train strong DA samples. Meanwhile, a group-based model aggregation is proposed to achieve a trade-off between supervised and unsupervised learning.
@@ II-C Semi-Asynchronous Federated Learning
Synchronous federated learning has become a natural choice, however, due to the device heterogeneity, the server has to wait for the slowest client in the synchronous protocol, which leads to low round efficiency, slow convergence speed, and low client utilization. To alleviate the above issues, asynchronous federated learning is proposed, whose essential idea is to perform aggregation once the server receives the model parameter uploaded by one client, without waiting for other clients, while allowing clients to use the stale global model for local training [23]. Stale control has been shown to be the key to guaranteeing model convergence.
@@
Wu et al. [17] allow some lag tolerance to leverage the knowledge of clients, while forcing to synchronize seriously outdated clients to prevent them from poisoning the global model. Ma et al. [18] conduct a theoretical analysis of the quantitative relationship between the convergence boundary of federated learning and different factors, such as the number of participating clients in each round, the degree of data non-IID distribution, and the heterogeneity of clients. Meanwhile, an adaptive learning rate based on client participation is designed to improve training accuracy. Chen et al. [31] propose a temporally weighted aggregation function, i.e., the most recently updated model should have a higher weight in the aggregation. In addition, the parameters of each layer of the CNN model are further analyzed to reduce the communication overhead. Chai et al. [32] stratify the clients according to the training speed, using the synchronous aggregation within the layer and asynchronous update between layers. Meanwhile, a lossy compression technique based on polyline encoding is adopted to reduce the communication cost.
@@ III-A Federated Learning
Federated Learning is first proposed in [7], whose idea is to allow users to cooperatively train a global model in order to profit from data of peer entities without sharing local raw data, i.e., ensuring data local privatization. A traditional federated learning system consists of a centralized server $S$ and multiple decentralized clients (ranging from several to thousands) $C_{i},1\leq i\leq M$. Clients perform local training based on their private data, and cooperatively train a global model under the coordination and command of the server, which acts as an aggregator. The procedure of federated learning is outlined as follows:
@@ III-A2 Local Training
where $\omega_{i}^{r}$ is the model parameter of client $C_{i}$ at round $r$, and $f(\omega_{i}^{r},x_{j},y_{j})$ is the loss function, which is used to measure the loss deviation between the prediction and the ground-truth of the data sample under the current model parameter $\omega_{i}^{r}$, such as cross-entropy loss [33].
Many optimization methods such as stochastic gradient descent (SGD) [34] or Adam [35] are used to find the optimal solution. The model parameter $\omega_{i}^{r}$ is updated locally by
@@ III-A3 Global Aggregation
The server $S$ performs the aggregation operation using a specific aggregation function, such as FedAvg [7], after receiving the local updated parameters of all selected clients to obtain the global model parameter for the next round of training. The aggregation function of FedAvg can be represented as
@@ III-B2 Security Gateways
Since IoT devices are moderately reliable and have limited resources, in this paper, the entity deploying and holding the anomaly detection component is not the IoT devices that need to be protected, but more powerful and reliable security gateways. Security gateways have sufficient resources to use effective approaches, such as Intel SGX [36] or remote attestation techniques [37], to ensure that they are not compromised. Security gateways act as the local access gateways, monitor the communication, network traffic or other data of the connected IoT devices, and use the trained model for anomaly detection. Meanwhile, they connect to the security service provider via the Internet or cellular, as clients in the federated learning system, uploading and downloading the models under the coordination of the security server provider.
@@ IV-B Federated Semi-Supervised Learning
In addition, the other works are based on the assumption that “IoT devices are not compromised in the model training stage, leaving sufficient time to learn models of benign behaviors”. Unfortunately, this assumption is not reasonable for vulnerable IoT devices with limited energy and computing resources. Once the devices are attacked and compromised during the training phase, malicious training data which is mistaken as benign will greatly reduce the accuracy of the model, whose effect is similar to data poisoning (label-flipping) attacks [38].
@@ IV-B1 Unsupervised Learning at Clients
Due to the lack of professional knowledge and the huge scale of private data, client-side data samples are usually not attached with any labels. Therefore, in this paper, we adopt a pseudo-labeling-based [39] training method to achieve unsupervised training on the unlabeled dataset at the clients. The loss function is defined as follows:
@@
where $\bar{y}=p(\omega_{i}^{r},x_{j})$, which means the prediction of the current model parameter $\omega_{i}^{r}$ on the sample $x_{j}$, $l(\ )$ is the cross-entropy loss function, $\operatorname{sgn}(\ )$ is the indicator function, and $\theta$ is the threshold hyperparameter, which is used to determine which samples with high confidence can be attached with pseudo-labels and used to help train the model.
@@ IV-C Semi-Asynchronous Federated Learning
Although synchronous FL has been widely used for anomaly detection in IoT networks, several limitations stand out as follows: 1) Low client utilization. Clients participating in each round of training are pre-specified and randomly selected. As a result, many capable clients remain idle even when they are ready and willing to participate in model training [17]. 2) Low round efficiency. Before aggregation at each round, the server has to wait for all selected clients to finish local training or reach the timeout threshold, however, due to the high heterogeneity of IoT devices, the time for devices to perform local updates varies greatly, resulting in low round efficiency. 3) Waste of training progress. Selected clients may not complete local training in time, and their progress may be wasted as clients participating in the next round of training are reassigned and their local models are forced to be overwritten by the global model.
Asynchronous FL is proposed to address the above shortcomings, but it also introduces the following new problems: 1) Significant computational and communication cost: In asynchronous FL, the server starts the global update upon receiving a model parameter from an arbitrary client, which results in significant resource consumption. 2) Bias in client contributions: due to the high heterogeneity of devices, clients with faster computing will frequently participate in the global update, resulting in a greater impact on the trained global model. 3) Poisoning of staleness: Some local models with large staleness can poison the global model, resulting in lower training accuracy, especially on non-IID data [40].

### v3

@@ Abstract
publicationid: pubid: 0000–0000/00\$00.00 © 2021 IEEE
@@ I Introduction
Over the past few decades, with the development of sensors and wireless communication technologies, Internet of Things (IoT) networks have become a popular architecture to support many modern applications or services, such as smart homes, smart healthcare, smart cities, smart grid, and smart transportation, among many others [MinFGCS]. However, simultaneously with the massive deployment of IoT networks, the attack surface where IoT devices may be hacked and exploited has grown in recent years, such as DDoS attack, Side-channel attack or Bot attack [IoTAttack1], which seriously hinders the development of IoT networks [IoTAttack]. For this reason, effective countermeasures need to be developed to detect infected IoT devices in networks, in order to strengthen and ensure the security of IoT networks.
In recent years, artificial intelligence (AI) technologies, such as machine learning (ML) and deep learning (DL), have been widely used for attack detection in IoT networks due to their superior performance [AIIoT]. Depending on the detection location, traditional AI-based schemes can be categorized into centralized and distributed detection [IoTSSuvery]. In centralized detection, IoT devices send their local data to a central server for detection. This inevitably introduces expensive communication costs, serious privacy leakage, and poor scalability, although it can achieve high detection accuracy. Distributed detection, known also as on-device detection, is performed locally on devices based on their private data. Nonetheless, it is limited by the lack of interaction with and profit from other devices, which leads to the problem of isolated data islands, resulting in low detection accuracy [COIoT].
To address the aforementioned limitations, Federated Learning (FL) [FedAvg] comes into being, in which multiple devices (i.e., clients) cooperatively train a global model under the coordination of a central entity (i.e., the server) without the leakage of local data. At each round of FL, clients perform the local training and upload the model parameters instead of raw data, and then the server is responsible for aggregating and distributing the parameters to share the knowledge in the data of various clients, thus achieving a trade-off between privacy, accuracy, communication cost and latency. The FL procedure continues until the model convergence or after a specified number of rounds. Many FL-based approaches for anomaly detection in IoT networks have been proposed, and Table I summarizes some classic and state-of-the-art works [DIoT, NBaIoTCN, ToNIoTCN, DADIoT, AFLIoT]. Based on the comprehensive analysis, we find that the existing researches are still subject to the following challenges:
@@
Nguyen et al. [DIoT] Full No Benign Synchronous None
Rey et al. [NBaIoTCN] Full Yes Yes/Benign Synchronous None
Campos et al. [ToNIoTCN] Full Yes Yes Synchronous None
Liu et al. [DADIoT] Full No Yes Synchronous Gradient Compression
Tian et al. [AFLIoT] One Yes Benign Asynchronous Gradient Compensation
@@
  • Lack of Labeled Data: A common limitation of existing works is that they only consider the supervised learning scenarios, where the local data is fully annotated with ground truths. It is not realistic because users and end devices do not have the ability and corresponding expertise to label the data [FedMatch]. Furthermore, some works [DIoT, AFLIoT] argue their schemes are unsupervised, but in fact, they are based on the assumption that all data in the model training phase is benign, which is essentially supervised. At the same time, this assumption is unreasonable because we can not guarantee it in real networks, and once the devices are compromised during the training phase, the malicious samples will seriously degrade the performance [Backdoor].
  • Device Heterogeneity: Although synchronous FL [DIoT, IoTSuvery1, IoTFL1] is a natural choice, due to the vast differences in computing capacities, communication resources and data volumes, the required time for different devices to perform local updates and receive/upload models varies greatly [SAFA]. Therefore, due to the device heterogeneity, the server has to wait for the slowest client in synchronous FL, which leads to unnecessary waiting time and low round efficiency. Although asynchronous FL can deal with the above problems, it requires frequent model updates, resulting in an expensive consumption of communication and computing resources. The server receives numerous local updates sent by clients, which can overwhelm the server and provide little benefit to the model convergence [FedSA].
  • Non-IID Data: In IoT networks, some devices may function normally, while others may suffer from multiple attacks. As a result, the local data of devices can not be regarded as a uniform distribution of the overall sample. In other words, the data of different clients is often not independent-and-identically-distributed (Non-IID) [NonIID2]. It has been proved that in the presence of non-IID data, the performance of FL degrades significantly in terms of the convergence rate and model accuracy [NonIID1].
  • Significant Communication Cost: Although FL does not need to upload local raw data to the server, frequent updates of model parameters (up to tens or even hundreds of rounds) are still a huge communication overhead for resource-constrained IoT networks. Meanwhile, the server may become a communication performance bottleneck since it needs to frequently distribute/receive the models to/from clients [CEFL1].
@@
  • We implement extensive experiments on the CIC-IDS 2017 [CIC-IDS] dataset. The experimental results show that the performance of FedS3A is far superior to the state-of-the-art detection methods [DIoT, NBaIoTCN, ToNIoTCN, DADIoT, AFLIoT] based on federated learning [FedAvg, AFL] in terms of the model performance, communication cost and round efficiency. It can achieve greater than 98% accuracy even when the data is non-IID and reduce higher than 50% communication cost. Furthermore, we conduct ablation experiments on FedS3A and evaluate the impact of different hyperparameters and functions to demonstrate its efficiency and effectiveness.
@@ II-A Federated Learning for Anomaly Detection in IoT Networks
Federated Learning has been widely used for anomaly detection in IoT networks due to cooperative learning and privacy preservation. DIoT [DIoT] is the first attempt to apply the federated learning approach for anomaly detection in IoT networks. It assumes that IoT devices are not compromised during the model training phase, leaving enough time to utilize the Gate Recurrent Unit (GRU) model to learn benign behavior patterns. Abnormalities are detected by deviations from benign behaviors. Liu et al. [DADIoT] use Convolutional Neural Networks (CNN) to extract fine-grained features, and then the trained Long Short-Term Memory (LSTM) model is utilized for anomaly detection. In addition, in order to improve communication efficiency, a top-k selection-based gradient compression scheme is proposed to reduce the communication cost.
Rey et al. [NBaIoTCN] studies the performance of federated learning in two scenarios based on the N-BaIoT [N-BaIoT] dataset. One is that the client-side data is all labeled, and MultiLayer Perceptron (MLP) is used to perform supervised learning for classification, while the other is that clients only have benign data and utilize AutoEncoder (AE) for anomaly detection training. Based on the CIC-ToN-IoT [CIC-ToN-IoT] dataset, Campos et al. [ToNIoTCN] define three scenarios, namely basic, balanced, and mixed scenarios, according to the distribution of the client-side dataset. Meanwhile, a novel aggregation function called Fed+ is proposed, which is more suitable for non-IID and skewed training data, and Multiple Linear Regression (MLR) is used for model training. Tian et al. [AFLIoT] focus on the gradient divergence and delay problem caused by the IoT device heterogeneity and non-IID data in federated learning. A method based on Taylor expansion is proposed to compensate for the inconsistency.
@@ II-B Federated Semi-Supervised Learning
Semi-supervised learning [Fixmatch] has attracted increasing attention in the past few years. It aims to learn a model under massive unlabeled data with a little supervised information. Therefore, in the context of federated learning, a novel and promising paradigm called federated semi-supervised learning (FSSL) is proposed and quickly becomes a research hotspot [lu2023framework]. There are two essential scenarios according to the location of the labeled data. The first is the conventional scenario where clients have both labeled and unlabeled data (labels-at-client), and the second is the disjoint scenario where the labeled data is only available at the server (labels-at-server) [FedMatch].
Zhu et al. [GPSFSSL] combine the CNN-GRU model with the pseudo-labeling technique for FSSL to identify the travel model. Meanwhile, a group-based aggregation function and a flip-based data augmentation (DA) scheme are proposed to improve the performance and convergence speed of the trained model. Nandury et al. [DLFSSL] analyze the influence of the update diversity of model parameters on model convergence, and propose diversity scaling to compensate for the effect of inter-client update dissimilarity. Zhang et al. [Benchmarking] propose a metric to measure the non-IID distribution of inter-client data. Pseudo-labels generated by weak DA samples are used to train strong DA samples. Meanwhile, a group-based model aggregation is proposed to achieve a trade-off between supervised and unsupervised learning.
@@ II-C Semi-Asynchronous Federated Learning
Synchronous federated learning has become a natural choice, however, due to the device heterogeneity, the server has to wait for the slowest client in the synchronous protocol, which leads to low round efficiency, slow convergence speed, and low client utilization. To alleviate the above issues, asynchronous federated learning is proposed, whose essential idea is to perform aggregation once the server receives the model parameter uploaded by one client, without waiting for other clients, while allowing clients to use the stale global model for local training [AFL]. Stale control has been shown to be the key to guaranteeing model convergence.
@@
Wu et al. [SAFA] allow some lag tolerance to leverage the knowledge of clients, while forcing to synchronize seriously outdated clients to prevent them from poisoning the global model. Ma et al. [FedSA] conduct a theoretical analysis of the quantitative relationship between the convergence boundary of federated learning and different factors, such as the number of participating clients in each round, the degree of data non-IID distribution, and the heterogeneity of clients. Meanwhile, an adaptive learning rate based on client participation is designed to improve training accuracy. Chen et al. [AFL1] propose a temporally weighted aggregation function, i.e., the most recently updated model should have a higher weight in the aggregation. In addition, the parameters of each layer of the CNN model are further analyzed to reduce the communication overhead. Chai et al. [FedAT] stratify the clients according to the training speed, using the synchronous aggregation within the layer and asynchronous update between layers. Meanwhile, a lossy compression technique based on polyline encoding is adopted to reduce the communication cost.
@@ III-A Federated Learning
Federated Learning is first proposed in [FedAvg], whose idea is to allow users to cooperatively train a global model in order to profit from data of peer entities without sharing local raw data, i.e., ensuring data local privatization. A traditional federated learning system consists of a centralized server $S$ and multiple decentralized clients (ranging from several to thousands) $C_{i},1\leq i\leq M$. Clients perform local training based on their private data, and cooperatively train a global model under the coordination and command of the server, which acts as an aggregator. The procedure of federated learning is outlined as follows:
@@ III-A2 Local Training
where $\omega_{i}^{r}$ is the model parameter of client $C_{i}$ at round $r$, and $f(\omega_{i}^{r},x_{j},y_{j})$ is the loss function, which is used to measure the loss deviation between the prediction and the ground-truth of the data sample under the current model parameter $\omega_{i}^{r}$, such as cross-entropy loss [CEL].
Many optimization methods such as stochastic gradient descent (SGD) [SGD] or Adam [Adam] are used to find the optimal solution. The model parameter $\omega_{i}^{r}$ is updated locally by
@@ III-A3 Global Aggregation
The server $S$ performs the aggregation operation using a specific aggregation function, such as FedAvg [FedAvg], after receiving the local updated parameters of all selected clients to obtain the global model parameter for the next round of training. The aggregation function of FedAvg can be represented as
@@ III-B2 Security Gateways
Since IoT devices are moderately reliable and have limited resources, in this paper, the entity deploying and holding the anomaly detection component is not the IoT devices that need to be protected, but more powerful and reliable security gateways. Security gateways have sufficient resources to use effective approaches, such as Intel SGX [SGX] or remote attestation techniques [RemoteAttestation], to ensure that they are not compromised. Security gateways act as the local access gateways, monitor the communication, network traffic or other data of the connected IoT devices, and use the trained model for anomaly detection. Meanwhile, they connect to the security service provider via the Internet or cellular, as clients in the federated learning system, uploading and downloading the models under the coordination of the security server provider.
@@ IV-B Federated Semi-Supervised Learning
In addition, the other works are based on the assumption that “IoT devices are not compromised in the model training stage, leaving sufficient time to learn models of benign behaviors”. Unfortunately, this assumption is not reasonable for vulnerable IoT devices with limited energy and computing resources. Once the devices are attacked and compromised during the training phase, malicious training data which is mistaken as benign will greatly reduce the accuracy of the model, whose effect is similar to data poisoning (label-flipping) attacks [DataPoisoning].
@@ IV-B1 Unsupervised Learning at Clients
Due to the lack of professional knowledge and the huge scale of private data, client-side data samples are usually not attached with any labels. Therefore, in this paper, we adopt a pseudo-labeling-based [Pseudo] training method to achieve unsupervised training on the unlabeled dataset at the clients. The loss function is defined as follows:
@@
where $\bar{y}=p(\omega_{i}^{r},x_{j})$, which means the prediction of the current model parameter $\omega_{i}^{r}$ on the sample $x_{j}$, l()l(\ ) is the cross-entropy loss function, $\operatorname{sgn}(\ )$ is the indicator function, and $\theta$ is the threshold hyperparameter, which is used to determine which samples with high confidence can be attached with pseudo-labels and used to help train the model.
@@ IV-C Semi-Asynchronous Federated Learning
Although synchronous FL has been widely used for anomaly detection in IoT networks, several limitations stand out as follows: 1) Low client utilization. Clients participating in each round of training are pre-specified and randomly selected. As a result, many capable clients remain idle even when they are ready and willing to participate in model training [SAFA]. 2) Low round efficiency. Before aggregation at each round, the server has to wait for all selected clients to finish local training or reach the timeout threshold, however, due to the high heterogeneity of IoT devices, the time for devices to perform local updates varies greatly, resulting in low round efficiency. 3) Waste of training progress. Selected clients may not complete local training in time, and their progress may be wasted as clients participating in the next round of training are reassigned and their local models are forced to be overwritten by the global model.
Asynchronous FL is proposed to address the above shortcomings, but it also introduces the following new problems: 1) Significant computational and communication cost: In asynchronous FL, the server starts the global update upon receiving a model parameter from an arbitrary client, which results in significant resource consumption. 2) Bias in client contributions: due to the high heterogeneity of devices, clients with faster computing will frequently participate in the global update, resulting in a greater impact on the trained global model. 3) Poisoning of staleness: Some local models with large staleness can poison the global model, resulting in lower training accuracy, especially on non-IID data [StaleNon-IID].